OpenStudy (anonymous):

How do you write x^1/2 * y^1/3 * z^1/6 as a single radical expression?

OpenStudy (anonymous):

gotta get that common denominator :) \[\frac{1}{2} = \frac{3}{6}, \frac{1}{3} = \frac{2}{6}\] so now we have: \[x^{\frac{3}{6}}y^{\frac{2}{6}}z^{\frac{1}{6}} = \sqrt[6]{x^3y^2z}\]